Determine the solution of the following differential equation:

xxy '= 1 + y^2


1
Expert's answer
2021-10-17T15:12:49-0400
"xy '= 1 + y^2"

"\dfrac{dy}{1+y^2}=\dfrac{dx}{x}"

Integrate


"\int \dfrac{dy}{1+y^2}=\int\dfrac{dx}{x}"

"\tan^{-1}y=\ln(|x|)+C"

"y=\tan(\ln(|x|)+C)"